OpenStudy (anonymous):

A stone is dropped into a lake, causing circular waves whose radii increase at a rate of 30 cm/sec. a. At what rate is the area enclosed by a circular wave increasing when the radius is 50 cm. b. At what rate is the circumference of a circular wave increasing when the radius is 50 cm?

OpenStudy (gorv):

dr/dt=30 area=pi*r^2 rate of change of area d(area)/dt=pi*2*r*dr/dt=pi*2*r*dr/dt dr/dt=30 d(area)/dt=2*pi*30*r=60*pi*50=3000*pi

OpenStudy (gorv):

circumference=2*pi*r rate of change=dc/dt=2*pi*dr/dt=2*pi*30=60pi

OpenStudy (gorv):

@sarcos11